Good evening, I need assistance with this interest problem and would greatly appreciate your help. Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
A. $496.30
B. $512.40
C. $555.40
D. $574.90
E. $588.20
the answer is A $496.30 but I do not understand how they have got the solution to that problem. what I have gotten so far is that I multiplied 670 with .35 and the answer is 234.5 then I took the answer and subtracted it from 670 which the answer 435.5

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
-----------
670 = Cost + 0.35*Cost = 1.35*Cost
Cost = 670/1.35

Answer by stanbon(75887)

Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
----
Let his cost be "x":
Equation:
x + 0.35x = 670
1.35x = 670
x = 670/1.35
x = $496.30
=================
Cheers,
Stan H.
=================
A. $496.30
B. $512.40
C. $555.40
D. $574.90
E. $588.20
the answer is A $496.30